OpenStudy (anonymous):

three times a number plus 12 minus 5 times the same number is 22. what is the number?

OpenStudy (anonymous):

Three times a number plus 12 minus 5 times the same number is 22 3n + 12 - 5n = 22 Combine like terms. 3n + 12 - 5n = 22 2n + 12 = 22 2n = 10 n = 10/2 What is n?